# RAG-Enhanced Collaborative LLM Agents for Drug Discovery

### v2

@@ Abstract
Recent advances in large language models (LLMs) have shown great potential to accelerate drug discovery. However, the specialized nature of biochemical data often necessitates costly domain-specific fine-tuning, posing critical challenges. First, it hinders the application of more flexible general-purpose LLMs in cutting-edge drug discovery tasks. More importantly, it impedes the rapid integration of the vast amounts of scientific data continuously generated through experiments and research. To investigate these challenges, we propose CLADD, a retrieval-augmented generation (RAG)-empowered agentic system tailored to drug discovery tasks. Through the collaboration of multiple LLM agents, CLADD dynamically retrieves information from biomedical knowledge bases, contextualizes query molecules, and integrates relevant evidence to generate responses — all without the need for domain-specific fine-tuning. Crucially, we tackle key obstacles in applying RAG workflows to biochemical data, including data heterogeneity, ambiguity, and multi-source integration. We demonstrate the flexibility and effectiveness of this framework across a variety of drug discovery tasks, showing that it outperforms general-purpose and domain-specific LLMs as well as traditional deep learning approaches.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLM) have revolutionized the landscape of natural language processing, emerging as general-purpose foundation models with remarkable abilities across multiple domains (Achiam et al., 2023; Touvron et al., 2023). In particular, their application in biomolecular studies has recently gained significant interest, motivated by the potential to profoundly accelerate scientific innovation and drug discovery applications (Zhang et al., 2024; Pei et al., 2024; Chaves et al., 2024). LLMs provide novel ways to understand and reason about molecular data, building on the wealth of available scientific literature. Additionally, their reasoning and zero-shot abilities help overcome the limitations of task-specific deep learning models, streamlining data needs and improving human-AI collaboration (Fang et al., 2023; Yu et al., 2024).
@@
However, given the inherent complexity and specialized nature of the field, recent works emphasize the importance of domain-specific fine-tuning to boost tasks such as molecular captioning, property prediction, or binding affinity prediction (Fang et al., 2023; Chaves et al., 2024; Yu et al., 2024; Edwards et al., 2024). Consequently, rather than employing readily available general-purpose LLMs, most efforts in drug discovery have focused on fine-tuning LLMs using biochemical annotations or instruction-tuning datasets.
@@
While promising, solely relying on these approaches poses significant challenges that can limit applications. On one hand, the rapid emergence of new LLM architectures and techniques (Minaee et al., 2024; Zhao et al., 2023b) complicates maintaining domain-specific models obtained through expensive fine-tuning. More importantly, drug discovery applications often require promptly incorporating new insights as they become available, for example, as a result of new experiments or through the scientific literature—a process exacerbated by the automation of experimental workflows (Tom et al., 2024). In addition to being impractical, regular rounds of fine-tuning to keep LLMs up-to-date with the latest scientific advances also introduce challenges such as catastrophic forgetting (Luo et al., 2023).
@@
From this perspective, retrieval-augmented generation (RAG) methods offer a promising solution that enables dynamic adaptation of the model’s knowledge without the need for expensive fine-tuning (Gao et al., 2023; Fan et al., 2024). However, applying this paradigm in the drug discovery domain presents important obstacles. First, retrieving relevant knowledge is difficult due to the limited domain expertise of general-purpose LLMs, combined with the vastness of the chemical space (Bohacek et al., 1996) that renders exact retrieval suboptimal. Second, biochemical data is extremely heterogeneous, spanning diverse modalities such as molecules, proteins, diseases, and complex relationships between them (Wang et al., 2023). Such data can also exist across multiple sources, introducing challenges in factual integration (Harris, 2023). Finally, the available information is not necessarily relevant to the query, as it may be too general, ambiguous, or partial (Vamathevan et al., 2019).
@@
In this study, we tackle these challenges by introducing a Collaborative framework of LLM Agents for Drug Discovery (CLADD). We assume a general setting where external knowledge is available as expert annotations associated with molecules or as knowledge graphs that flexibly represent diverse biochemical entities and their relationships. CLADD is powered by general-purpose LLMs, while also integrating domain-specific LMs, when necessary, to improve molecular understanding. Notably, external knowledge can be updated dynamically without LLM fine-tuning.
@@
The multi-agent collaborative framework enables each agent to specialize in a specific data source and/or role based on their team, offering a modular solution that can improve overall information processing (Chan et al., 2024). In particular, CLADD includes a Planning Team to determine relevant data sources, a Knowledge Graph Team to retrieve external heterogeneous information in the knowledge graph and summarize it, also through a novel anchoring approach to retrieve related information when the query molecule is not present in the knowledge base, and a Molecule Understanding Team, which analyzes the query molecule based on its structure along with summaries of external data and tools. The flexibility of the framework enables CLADD to address a wide range of tasks for drug discovery, including zero-shot settings, while also improving interpretability through the transparent interaction of its agents.
@@
  • We present CLADD, a multi-agent framework for RAG-based question-answering in drug discovery applications. The framework leverages generalist LLMs and dynamically integrates external biochemical data from multiple sources without requiring fine-tuning.
  • We demonstrate the flexibility of the framework by tackling diverse applications, including property-specific molecular captioning, drug-target prediction, and molecular toxicity prediction.
  • We provide comprehensive experimental results showcasing the effectiveness of CLADD compared to general-purpose and domain-specific LLMs, as well as standard deep learning approaches. A further appeal of CLADD is its flexibility and explainability, improving the interaction between scientists and AI.
Figure 1: Overview of CLADD. (a) General overview. Given a question about a molecule (SMILES representation), the Planning Team first evaluates the relevant data sources and models. If the knowledge graph is considered relevant, a dedicated report will be created by the Knowledge Graph Team. The Molecule Understanding Team creates a report on the molecule based on the relevant annotation databases and tools. The generated reports are integrated by the Prediction Agent and used to produce the final answer. (b) Detailed overview of the Planning Team. The Molecular Annotation Planner decides whether to use additional captioning tools to complement the information associated with the molecule. (c) Detailed overview of the Knowledge Graph Team. The Drug Relation and Biological Relation Agents retrieve information from the knowledge graph based on the anchor drug, the related drugs, and the set of relations between the anchor and related drugs (cfr. Section 3.2.2).
@@ 2 Related Work
LLMs for Molecules. Leveraging the extensive body of literature and string-based molecular representations such as SMILES, language models (LMs) have been successfully applied to molecular sciences. Inspired by the masked language modeling approach used in BERT training (Devlin et al., 2018), KV-PLM (Zeng et al., 2022) introduces a method to train LMs by reconstructing masked SMILES and textual data. Similarly, MolT5 (Edwards et al., 2022) adopts the “replace corrupted spans” objective (Raffel et al., 2020) for pre-training on both SMILES strings and textual data, followed by fine-tuning for downstream tasks such as molecule captioning and generation. Building on this foundation, Pei et al. (2023) and Christofidellis et al. (2023) extend MolT5 with additional pre-training tasks, including protein FASTA reconstruction and chemical reaction prediction. Furthermore, GIMLET (Zhao et al., 2023a), Mol-Instructions (Fang et al., 2023), and MolecularGPT (Liu et al., 2024b) adopt instruction tuning (Zhang et al., 2023) to improve generalization across a wide range of molecular tasks. While these approaches demonstrate enhanced versatility, they still rely on expensive fine-tuning processes to enable molecule-specific tasks or to incorporate new data.
@@
LLM Agents for Science. An LLM agent is a system that leverages LLMs to interact with users or other systems, perform tasks, and make decisions autonomously (Wang et al., 2024a). Recently, LLM agents have attracted significant interest in scientific applications and biomedical discovery (Gao et al., 2024), with applications including literature search (Lála et al., 2023), experiment design (Roohani et al., 2024), and hypothesis generation (Wang et al., 2024b), among others. In particular, agents focusing on drug discovery applications have emerged. Systems like ChemCrow (Bran et al., 2023), CACTUS (McNaughton et al., 2024), and Coscientist (Boiko et al., 2023) focus on automating cheminformatics tasks and experiments, streamlining computational and experimental pipelines. Other works leverage agent-based orchestration of tools and data to accelerate specific aspects of scientific workflows, such as search (ODonoghue et al., 2023) or design (Ghafarollahi & Buehler, 2024). In contrast to existing works, we investigate an agent-based framework that can effectively incorporate external knowledge to improve general molecular QA. This could be used either independently or as part of a larger system for automated drug discovery (Tom et al., 2024).
@@
Multi-Agent Collaborations for Drug Discovery. Only a limited number of studies have explored multi-agent frameworks in the context of drug discovery. DrugAgent (Inoue et al., 2024) introduces a multi-agent framework integrating multiple external data sources but is limited to predicting drug-target interaction scores. Another study with the same name employs an agentic framework for automating machine learning programming for drug discovery tasks (Liu et al., 2024a). In contrast, our work seeks to tackle a diverse array of drug discovery tasks, enabling applicability across a wide variety of use cases. We provide additional related works on the knowledge graph in Appendix A.
@@ 3 Methodology
3.1 Problem Setup
[…]
We assume access to two types of external databases: (1) molecular annotation databases $\mathcal{C}$, which contain textual annotation about molecules (for example, detailing their functions and properties) and (2) knowledge graphs (KGs) connecting molecules to other biomedical entities. In particular, a KG $\mathcal{G}$ is composed of a set of entities $\mathcal{E}$ and a set of relations $\mathcal{R}$ connecting them. KG can include various types of entities, such as drugs, proteins, and diseases. In this paper, we only assume that molecule (or drug) entities are present in KG, while any other types of entities can exist.
Additionally, we assume access to pre-trained molecular captioning models that can be used as external tools to complement the external databases. In general, any predictive model on molecules can be considered a captioning model (Edwards et al., 2022; Pei et al., 2023), given that its output can be simply represented as text.
@@ 3.2 CLADD
Here, we introduce CLADD, a multi-agent framework for general molecular question-answering that supports multiple drug discovery tasks. Each agent is implemented by an off-the-shelf LLM prompted to elicit a particular behavior. Our framework is composed of three teams, each composed of several agents: the Planning Team, which identifies the most appropriate data sources and overall strategy given the task and the query molecule (Section 3.2.1); the Knowledge Graph (KG) Team, which retrieves relevant contextual information about the molecule from available KG databases (Section 3.2.2); and the Molecular Understanding (MU) Team, which retrieves and integrates information from molecular annotation databases and external tools for molecule description (Section 3.2.3); Finally, the Prediction Agent integrates the findings from the MU and KG teams to generate the final answer. In the following sub-sections, we describe each team in detail. The overall framework is depicted in Figure 1.
3.2.1 Planning Team
@@
The Planning Team assesses prior knowledge for a given query molecule. The team separately assesses the relevance of the molecular annotations and the knowledge graph using a MolAnn Planner agent and a KG Planner agent.
@@
Molecule Annotation (MolAnn) Planner. This agent first retrieves annotations for the query molecule, $c_{q}$, from the annotation database $\mathcal{C}$. While these annotations can provide valuable biochemical knowledge (Yu et al., 2024), they are often sparse and may lack sufficient details due to the vastness of the chemical space (Lee et al., 2024). Given the vastness of the chemical space, it is also not uncommon for molecules to be completely absent from databases.
@@
To this end, the MolAnn Planner determines whether the retrieved annotations provide enough information for subsequent analyses. Specifically, given a query molecule $g_{q}$ and retrieved annotations $c_{q}$, the agent is invoked as follows:
$\small o_{\text{MAP}}=\text{MolAnn Planner}(g_{q},c_{q}).$ (1)
@@
$o_{\text{MAP}}$ is a Boolean indicating whether annotations should be complemented with additional information from tools.
@@
Knowledge Graph (KG) Planner. In parallel to analyzing the available description for the query molecule, we analyze the relevance of the contextual information present in the KG. Unlike previous works in general QA tasks, which primarily rely on identifying exact entity matches within the KG (Baek et al., 2023), the vast chemical search space and the limited coverage of existing knowledge bases hinder such approaches.
@@
To address this challenge, we propose leveraging the knowledge of drugs that are structurally similar to the query drug, building upon the well-established biochemical principle that structurally similar molecules often exhibit related biological activity (Martin et al., 2002). Specifically, we define the anchor drug $g_{a}$ as the entity drug that maximizes the cosine similarity between its embedding and that of the query molecule, among the set of all molecules in the KG ($g_{\mathcal{G}}$):
$\small g_{a}=\underset{g\in g_{\mathcal{G}}}{\text{argmax}}\,\frac{\textit{emb}(g_{q})\cdot\textit{emb}(g)}{\|\textit{emb}(g_{q})\|\|\textit{emb}(g)\|},$ (2)
where emb is a graph neural network (GNN) pre-trained with 3D geometry (Liu et al., 2021) that outputs structure-aware molecular embeddings.
@@
$\small o_{\text{KGP}}=\text{KG Planner}(g_{q},g_{a},s_{q,a}),$ (3)
@@
3.2.2 Knowledge Graph Team
@@
This team aims to provide relevant contextual information about the query molecule by leveraging the KG, and it is only called if $o_{\text{KGP}}=\text{TRUE}$. It consists of the Drug Relation (DrugRel) Agent and the Biological Relation (BioRel) Agent, both of which generate reports on the query molecule based on different aspects of the KG. Specifically, the DrugRel Agent focuses on related drug entities within the KG, primarily leveraging its internal knowledge, whereas the BioRel Agent focuses on summarizing and assessing biological relationships between drugs present in the KG.
@@
Related Drugs Retrieval. The typical approach to leveraging a KG for QA tasks involves identifying multiple entities in the query and extracting the subgraph that encompasses those entities (Baek et al., 2023; Wen et al., 2023). However, in molecular understanding for applications related to drug discovery tasks, the question often involves only a single entity, i.e., the query molecule $g_{q}$, making it challenging to identify information in the KG relevant to the task.
@@
Here, we introduce a novel approach for extracting relevant information for the query molecule $g_{q}$ by utilizing the retrieved anchor drug $g_{a}$, which exhibits high structural similarity to the query molecule. In particular, while the drug entities in the KG $\mathcal{G}$ are mainly connected to other types of biological entities (e.g., proteins, diseases), we can infer relationships among drugs by considering the biological entities they share. For example, we can determine the relatedness of the drugs Trastuzumab and Lapatinib by observing their connectivity to the protein HER2 in the KG, as both drugs specifically target and inhibit HER2 to treat HER2-positive breast cancer (De Azambuja et al., 2014). Therefore, to identify relevant related drugs, we first compute the 2-hop paths connecting the anchor drug $g_{a}$ to other drugs $g_{\mathcal{G}}^{i}$ in the KG $\mathcal{G}$, i.e., $(g_{a},r_{a\rightarrow e},e,r_{i\rightarrow e},g_{\mathcal{G}}^{i})$, where $r\in\mathcal{R}$, $e\in\mathcal{E}$, and $i$ denotes the index of the other drug. Then, we select the top-$k$ related drugs, denoted as $g_{r^{1}},\ldots,g_{r^{k}}$, corresponding to the molecules that have the greatest number of 2-hop paths to the anchor drug. Note that while the anchor drug $g_{a}$ is selected based on its structural similarity to the query molecule $g_{q}$, these reference drugs are biologically related to $g_{a}$, reflecting the relationships captured within the KG.
@@
Drug Relation (DrugRel) Agent. The DrugRel Agent generates a report on the query molecule, contextualizing it in relation to relevant drugs present in the knowledge base for the specific task instruction. Given a query molecule $g_{q}$, its anchor drug $g_{a}$, and the set of related drugs $g_{r^{1}},\ldots,g_{r^{k}}$, the DrugRel Agent is defined as follows:
$\small o_{\text{DRA}}=\text{DrugRel Agent}~{}(g_{q},g_{a},g_{r^{1}},\ldots,g_{r^{k}},\mathcal{T},\mathcal{I}),$ (4)
@@
where $\mathcal{T}=\{s_{q,a},s_{q,r^{1}},\ldots,s_{q,r^{k}}\}$ is the set of Tanimoto similarities between the query molecule and the retrieved drugs, and $\mathcal{I}$ is the task instruction. This allows the agent to leverage its internal knowledge about related drugs while effectively assessing the relatedness of the information to the target molecule based on the Tanimoto similarity.
@@
Biological Relation (BioRel) Agent. On the other hand, drugs related to the query molecule in the KG may exhibit limited structural similarity, underscoring the importance of utilizing additional relevant information, such as shared toxicity profiles or interactions with the same target, rather than relying solely on structural resemblance. Therefore, the BioRel Agent summarizes how the anchor drug and the related drugs are biologically related, integrating additional biochemical information present in the KG. Specifically, given an anchor drug $g_{a}$, a set of reference drugs $g_{r^{1}},\ldots,g_{r^{k}}$, the collection of all 2-hop paths $\mathcal{P}$ linking the anchor drug to the reference drugs, and the instruction $\mathcal{I}$, the agent generates the report as follows:
$\small o_{\text{BRA}}=\text{BioRel Agent}(\mathcal{P},\mathcal{I},g_{q},g_{a},s_{q,a}).\vspace{-5pt}$ (5)

### v3

@@ Abstract
Recent advances in large language models (LLMs) have shown great potential to accelerate drug discovery. However, the specialized nature of biochemical data often necessitates costly domain-specific fine-tuning, posing major challenges. First, it hinders the application of more flexible general-purpose LLMs for cutting-edge drug discovery tasks. More importantly, it limits the rapid integration of the vast amounts of scientific data continuously generated through experiments and research. Compounding these challenges is the fact that real-world scientific questions are typically complex and open-ended, requiring reasoning beyond pattern matching or static knowledge retrieval. To address these challenges, we propose CLADD, a retrieval-augmented generation (RAG)-empowered agentic system tailored to drug discovery tasks. Through the collaboration of multiple LLM agents, CLADD dynamically retrieves information from biomedical knowledge bases, contextualizes query molecules, and integrates relevant evidence to generate responses — all without the need for domain-specific fine-tuning. Crucially, we tackle key obstacles in applying RAG workflows to biochemical data, including data heterogeneity, ambiguity, and multi-source integration. We demonstrate the flexibility and effectiveness of this framework across a variety of drug discovery tasks, showing that it outperforms general-purpose and domain-specific LLMs as well as traditional deep learning approaches. Our code is publicly available at https://github.com/Genentech/CLADD.
@@ 1 Introduction
Large language models (LLM) have revolutionized the landscape of natural language processing, emerging as general-purpose foundation models with remarkable abilities across multiple domains (achiam2023gpt; touvron2023llama). In particular, their application in biomolecular studies has recently gained significant interest, motivated by the potential to profoundly accelerate scientific innovation and drug discovery applications (zhang2024scientific; pei2024leveraging). LLMs provide novel ways to understand and reason about molecular data, building on the wealth of available scientific literature. Additionally, their reasoning and zero-shot abilities help overcome the limitations of task-specific deep learning models, streamlining data needs and improving human-AI collaboration (fang2023mol).
@@
However, given the inherent complexity and specialized nature of the field, recent works emphasize the importance of domain-specific fine-tuning to boost tasks such as molecular captioning, property prediction, or binding affinity prediction (fang2023mol; chaves2024tx; yu2024llasmol; edwards2024molcap). Consequently, rather than employing readily available general-purpose LLMs, most efforts in drug discovery have focused on fine-tuning LLMs using biochemical annotations or instruction-tuning datasets.
@@
While promising, solely relying on these approaches poses significant challenges that can limit applications. On one hand, the rapid emergence of new LLM architectures and techniques (minaee2024large) complicates maintaining domain-specific models through expensive fine-tuning. More importantly, drug discovery applications often require promptly incorporating new insights as they become available, for example, through new experiments or the scientific literature. In addition to being impractical, regular rounds of fine-tuning also introduce challenges such as catastrophic forgetting (luo2023empirical), while not necessarily providing grounded answers (gekhman-etal-2024-fine). Above all, real-world drug discovery questions are inherently complex, open-ended, and context-dependent, spanning heterogeneous data types (ramos2025review). As a consequence, static LLMs—either general-purpose or fine‑tuned—may struggle to generalize to novel tasks or adapt to new evidence.
@@
From this perspective, retrieval-augmented generation (RAG) methods offer a promising direction that enables dynamic adaptation of the model’s knowledge without the need for continuous, expensive fine-tuning (gao2023retrieval; fan2024survey). However, applying this paradigm in the drug discovery domain presents important obstacles. First, retrieving relevant knowledge is difficult due to the limited domain expertise of general-purpose LLMs, combined with the vastness of the biochemical space (bohacek1996art) that renders exact retrieval ineffective. Second, biochemical data is extremely heterogeneous, spanning diverse modalities such as molecules, proteins, diseases, and complex relationships between them (wang2023scientific), which can also exist across multiple sources. Finally, many real-world tasks are open-ended and require the LLM to extrapolate beyond the available external knowledge, (which may also be ambiguous or partial (vamathevan2019applications)) while remaining grounded in it.
@@
In this study, we tackle these challenges by introducing a Collaborative framework of LLM Agents for Drug Discovery (CLADD). We assume a general setting where external knowledge is available as expert annotations associated with molecules or as knowledge graphs (KGs) that flexibly represent heterogeneous biochemical entities and their relationships. CLADD is powered by general-purpose LLMs, while also integrating domain-specific LLMs, to improve molecular understanding. Notably, external knowledge can be updated dynamically without LLM fine-tuning.
@@
The multi-agent collaborative framework enables each agent to specialize in a specific data source and/or role, offering a modular solution that can improve overall information processing (chan2024chateval). In particular, CLADD includes a Planning Team to determine relevant data sources, a Knowledge Graph Team to retrieve external heterogeneous information in the KG and summarize it, also through a novel anchoring approach to retrieve related information when the query molecule is not present in the knowledge base, and a Molecule Understanding Team, which analyzes the query molecule based on its structure, along with summaries of external data and tools. The flexibility of the framework enables CLADD to address a wide range of tasks for drug discovery, including zero-shot and open-ended settings, while also improving interpretability through the transparent interaction of its agents.
@@
  • We present CLADD, a multi-agent framework for RAG-based question-answering in drug discovery applications. The framework leverages generalist LLMs and dynamically integrates external biochemical data without fine-tuning, while addressing open-ended settings.
  • We demonstrate the flexibility of the framework by tackling diverse applications, including drug-target prediction, property-specific molecular captioning, and biological activity prediction tasks.
  • We provide comprehensive experimental results showcasing the effectiveness of CLADD compared to both general-purpose and domain-specific LLMs, as well as standard deep learning approaches.
Figure 1: Overview of CLADD.
@@ 2 Related Work
LLMs for Molecules. Leveraging the extensive body of literature and string-based molecular representations such as SMILES, language models (LMs) have been successfully applied to molecular sciences. Inspired by the masked language modeling approach used in BERT training (devlin2018bert), KV-PLM (zeng2022deep) introduces a method to train LMs by reconstructing masked SMILES and textual data. Similarly, MolT5 (edwards2022translation) adopts the “replace corrupted spans” objective (raffel2020exploring) for pre-training on both SMILES strings and textual data, followed by fine-tuning for downstream tasks such as molecule captioning and generation. Building on this foundation, pei2023biot5 and christofidellis2023unifying extend MolT5 with additional pre-training tasks, including protein FASTA reconstruction and chemical reaction prediction. Furthermore, GIMLET (zhao2023gimlet), Mol-Instructions (fang2023mol), and MolecularGPT (liu2024moleculargpt) adopt instruction tuning (zhang2023instruction) to improve generalization across a wide range of molecular tasks. While these approaches demonstrate enhanced versatility, they still rely on expensive fine-tuning processes to enable molecule-specific tasks or to incorporate new data.
@@
LLM Agents for Science. An LLM agent is a system that leverages LLMs to interact with users or other systems, perform tasks, and make decisions autonomously (wang2024survey). Recently, LLM agents have attracted significant interest in scientific applications and biomedical discovery (gao2024empowering), with applications including literature search (lala2023paperqa), experiment design (roohani2024biodiscoveryagent), and hypothesis generation (wang-etal-2024-scimon), among others. In particular, agents focusing on drug discovery applications have emerged. Systems like ChemCrow (bran2023chemcrow), CACTUS (mcnaughton2024cactus), and Coscientist (boiko2023autonomous) focus on automating cheminformatics tasks and experiments, streamlining computational and experimental pipelines. Other works leverage agent-based orchestration of tools and data to accelerate specific aspects of scientific workflows, such as search (odonoghue2023bioplanner) or design (ghafarollahi2024protagents). In contrast to existing works, we investigate an agent-based framework that can effectively incorporate external knowledge to improve open-ended and zero-shot molecular QA. This could be used either independently or as part of a larger system for automated drug discovery.
@@
Multi-Agent Collaborations for Drug Discovery. Only a limited number of studies have explored multi-agent frameworks in the context of drug discovery. DrugAgent (inoue2024drugagent) introduces a multi-agent framework integrating multiple external data sources, but is limited to predicting drug-target interaction. Another study with the same name employs an agentic framework for automating machine learning programming for drug discovery tasks (liu2024drugagent). In contrast, our work seeks to tackle a diverse array of drug discovery tasks, grounding the agent capabilities in external knowledge. We provide further related works on knowledge graph-based RAG in Appendix A.
@@ 3 Methodology
Problem Setup. Given a query molecule $g_{q}$ and a textual prompt describing a task of interest $\mathcal{I}$, we consider the general problem of generating a relevant response $\mathcal{A}_{g_{q}}$. For instance, given $g_{q}=\text{`C1=CC(=C(C=C1CCN)O)O '}$ and $\mathcal{I}=\text{`\emph{Predict liver toxicity}'}$, our model should be able to generate an answer stating that $\mathcal{A}_{g_{q}}=\text{`\emph{this molecule does not have liver toxicity concerns}'}$. Such a general QA setup can be flexibly adapted to multi-class classification, captioning, and set-based predictions.
We assume access to two types of external databases: (1) molecular annotation databases $\mathcal{C}$, which include textual annotation about molecules (for example, detailing their functions and properties), and (2) knowledge graphs (KGs) connecting molecules to other biomedical entities. In particular, a KG $\mathcal{G}$ is composed of a set of heterogeneous entities $\mathcal{E}$ (such as drugs, proteins, and diseases) and a set of relations $\mathcal{R}$ connecting them. In this paper, we only assume that molecule (or drug) entities are present in the KG, while any other types of entities can exist. Additionally, we assume access to pre-trained molecular captioning models that can be used as external tools to complement the external databases. In general, any predictive model on molecules can be considered a captioning model (edwards2022translation; pei2023biot5), given that its output can be simply represented as text.
@@ 3.1 CLADD
Here, we introduce CLADD, a multi-agent framework for general molecular question-answering that supports multiple drug discovery tasks. Each agent is implemented by an off-the-shelf LLM prompted to elicit a particular behavior. Our framework is composed of three teams, each composed of several agents: the Planning Team, which identifies the most appropriate data sources and overall strategy given the task and the query molecule (Section 3.1); the Knowledge Graph (KG) Team, which retrieves relevant contextual information about the molecule from available KG databases (Section 3.1); and the Molecular Understanding (MU) Team, which retrieves and integrates information from molecular annotation databases and external tools for molecule description (Section 3.1). Finally, the Prediction Agent integrates the findings from the MU and KG teams to generate the final answer. In the following sub-sections, we describe each team in detail. The overall framework is depicted in Figure 1.
Planning Team
@@
The Planning Team assesses the relevance of external knowledge for a given query molecule. The team separately assesses the molecular annotations database and the knowledge graph through the MolAnn Planner and the KG Planner agents, respectively.
@@
Molecule Annotation (MolAnn) Planner. This agent first retrieves annotations for the query molecule, $c_{q}$, from the annotation database $\mathcal{C}$. While these annotations can provide valuable biochemical knowledge (yu2024llasmol), they are often sparse, with many molecules entirely missing or lacking sufficient details due to the vastness of the chemical space (lee2024vision).
@@
To this end, the MolAnn Planner determines whether the retrieved annotations provide enough information for subsequent analyses. Specifically, given a query molecule $g_{q}$, retrieved annotations $c_{q}$, and the task instruction $\mathcal{I}$, the agent is invoked as follows:
$o_{\text{MAP}}=\text{MolAnn Planner}(g_{q},c_{q},\mathcal{I}).$ (1)
@@
$o_{\text{MAP}}$ indicates whether annotations should be complemented with additional information from tools.
@@
Knowledge Graph (KG) Planner. In parallel to analyzing the available description for the query molecule, we analyze the relevance of the contextual information present in the KG. While previous works on general QA tasks focus on identifying entities in the knowledge graph that exactly match those in the query (baek2023knowledge; jiang-etal-2023-structgpt), the vast chemical search space and the limited coverage of existing knowledge bases limit the effectiveness of such approaches in the field of drug discovery.
@@
To address this challenge, we propose leveraging the knowledge of drugs that are structurally similar to the query drug, building upon the well-established biochemical principle that structurally similar molecules often exhibit related biological activity (martin2002structurally). Specifically, we define the anchor drug $g_{a}$ as the entity drug with the maximum cosine similarity between its embedding and that of the query molecule, among the set of all molecules in the KG ($g_{\mathcal{G}}$), $g_{a}=\underset{g\in g_{\mathcal{G}}}{\text{argmax}}\,\frac{\textit{emb}(g_{q})\cdot\textit{emb}(g)}{\|\textit{emb}(g_{q})\|\|\textit{emb}(g)\|}$, where emb is a representation produced by a graph neural network (GNN) pre-trained with 3D geometry (liu2021pre), which outputs structure-aware molecular embeddings.
@@
$o_{\text{KGP}}=\text{KG Planner}(g_{q},g_{a},s_{q,a},\mathcal{I}),$ (2)
@@
Knowledge Graph Team
@@
This team aims to provide relevant contextual information about the query molecule by leveraging the KG, and it is only called if $o_{\text{KGP}}=\text{TRUE}$. It consists of the Drug Relation (DrugRel) Agent and the Biological Relation (BioRel) Agent, both of which generate reports on the query molecule based on different aspects of the KG. Specifically, the DrugRel Agent focuses on related drug entities within the KG, whereas the BioRel Agent focuses on summarizing and assessing contextual biological knowledge in the KG.
@@
Related Drugs Retrieval. The typical approach to leveraging a KG for QA tasks involves identifying multiple entities in the query and extracting the subgraph that encompasses those entities (wen2023mindmap). However, in molecular understanding for applications related to drug discovery tasks, the question often involves only a single entity, i.e., the query molecule $g_{q}$, making it challenging to identify information in the KG relevant to the task.
@@
Here, we introduce a novel approach for extracting relevant information for the query molecule $g_{q}$ by utilizing the retrieved anchor drug $g_{a}$, which exhibits high structural similarity to the query molecule. In particular, while the drug entities in the KG $\mathcal{G}$ are mainly connected to other types of biological entities (e.g., proteins, diseases), we can infer relationships among drugs by considering the biological entities they share. For example, we can determine the relatedness of the drugs Trastuzumab and Lapatinib by observing their connectivity to the protein HER2 in the KG, as both drugs specifically target and inhibit HER2 to treat HER2-positive breast cancer (de2014lapatinib). Therefore, to identify relevant related drugs, we first compute the 2-hop paths connecting the anchor drug $g_{a}$ to other drugs $g_{\mathcal{G}}^{i}$ in the KG $\mathcal{G}$, i.e., $(g_{a},r_{a\rightarrow e},e,r_{i\rightarrow e},g_{\mathcal{G}}^{i})$, where $r\in\mathcal{R}$, $e\in\mathcal{E}$, and $i$ denotes the index of the other drug. Then, we select the top-$k$ related drugs, denoted as $g_{r^{1}},\ldots,g_{r^{k}}$, corresponding to the molecules that have the greatest number of 2-hop paths to the anchor drug. Note that while the anchor drug $g_{a}$ is selected based on its structural similarity to the query molecule $g_{q}$, these reference drugs are semantically related to $g_{a}$, reflecting the relationships captured within the KG.
@@
Drug Relation (DrugRel) Agent. The DrugRel Agent generates a report on the query molecule, contextualizing it in relation to relevant drugs present in the knowledge base for the specific task instruction. Given a query molecule $g_{q}$, its anchor drug $g_{a}$, and the set of related drugs $g_{r^{1}},\ldots,g_{r^{k}}$, the DrugRel Agent generates a report as follows:
$o_{\text{DRA}}=\text{DrugRel Agent}~(g_{q},g_{a},g_{r^{1}},\ldots,g_{r^{k}},\mathcal{T},\mathcal{I}),$ (3)
@@
where $\mathcal{T}=\{s_{q,a},s_{q,r^{1}},\ldots,s_{q,r^{k}}\}$ is the set of Tanimoto similarities between the query molecule and the retrieved drugs. The agent leverages its internal knowledge about related drugs while effectively assessing the relatedness of the information to the target molecule based on the Tanimoto similarity.
@@
Biological Relation (BioRel) Agent. The BioRel Agent summarizes how the anchor drug and the related drugs are biologically related, integrating additional biochemical entities present in the KG, such as targets, indications, side effects, etc. Specifically, given an anchor drug $g_{a}$, a set of reference drugs $g_{r^{1}},\ldots,g_{r^{k}}$, the collection of all 2-hop paths $\mathcal{P}$ linking the anchor drug to the reference drugs, and the instruction $\mathcal{I}$, the agent generates the report as follows:
$o_{\text{BRA}}=\text{BioRel Agent}(\mathcal{P},\mathcal{I},g_{q},g_{a},s_{q,a}).\vskip-5.0pt$ (4)